OpenStudy (anonymous):

What is the simplified form of the expression 8√v^2 − √v^4 − √9v^2 + √16v^4 ? Answer Options: 15v^2 + 11v 11v^2 + 15v 5v^2 + 3v 3v^2 + 5v

OpenStudy (istim):

Sorry. I can't remember this stuff.

OpenStudy (anonymous):

thats ok I just thought I should put my question up and stop procrastinating in the chat area. I've been trying to figure out this particular question for a long time.

OpenStudy (istim):

If there's a point when you really must ask a a question, just private/direct message people, or call them by typing @name. For example. @IsTim

OpenStudy (anonymous):

@lucy4104

OpenStudy (istim):

@nincompoop

OpenStudy (istim):

Hey, could you draw out the equation? Or use the equation editor? That would help anyone trying to intepret what you are asking quicker.

OpenStudy (anonymous):

\[8\sqrt{v ^{2}} - \sqrt{v ^{4}} - \sqrt{9v ^{2}} + \sqrt{16v ^{4}}\] simplified

OpenStudy (tkhunny):

\(\sqrt{v^{4}} = v^{2}\) \(\sqrt{v^{2}} = |v|\)

OpenStudy (anonymous):

@tkhunny Okay so that would be the simplified term for the second term then? what about the fourth term? how should I simplify it?

OpenStudy (anonymous):